OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x – 2. What is f(x) • g(x)?

OpenStudy (anonymous):

f*g = (..computation of f..)*(computation of g)

OpenStudy (anonymous):

???

OpenStudy (anonymous):

i have no idea what that means

OpenStudy (anonymous):

Can you simply multiply the two functions: \[f(x) \cdot g(x) = (x + 7)(x-2) = ?\] Calculate this..

OpenStudy (anonymous):

is he asking for composition or product?